OpenStudy (babyslapmafro):

Please help me eliminate the following parameter in order to sketch the curve. x=cos(2t) y=sin(t)

OpenStudy (babyslapmafro):

yes that is correct....but how do i work with that?

OpenStudy (perl):

x= cos(2t) (given) = 1 - 2(sin(t))^2 (because trig identity) = 1 - 2*y^2 (substitute y=sin(t) )